OpenStudy (anonymous):

(A man is making a nut mixture of almounds worth $4 per pound and cashews worth $6 per pound. If there are 3 pounds less cashews than almonds, and the total mixture is worth $122, how many pounds of each is there)

OpenStudy (anonymous):

let x=almonds and y=cashews. now x= 3+y and 4x + 6y = 122...now solve simultaneously for x and y

OpenStudy (anonymous):

that is certainly one way. i would write \[4A+6(A-3)=122\] and solve for A